Is the function h(u,v)=cos⁡(u)+e^nu+√(sinh⁡(u^2 )+√(v+e^nu ))  , Where n is second digit of your arid number for example if your arid number is 19-arid-12345 then choose n=2. Is is continuous? In case of yes then find the points of continuity. Explain briefly


1
Expert's answer
2021-05-07T09:55:38-0400

Given,

"f(x)=x^2+2x"


"x^2+2x-y=0"


The solution of the above equation is-


"x=\dfrac{-2\pm\sqrt{4+4y}}{2}"


"x=-1\pm\sqrt{1+y}"


which is defined in R for "y\u2265\u22121." This can be used to separate the domain of f(x) into two intervals. For "x\in[1,∞" ) and "x\in(−∞,1]" , the function f(x) is not injective (and invertible). The domain that contains 2 is of course "[1,\u221e)."